OpenStudy (anonymous):

the montoyas need to travel 395 miles to get to their cousin's house. if they average 65mph , about how many hours will it take them to reach their destination ?

OpenStudy (anonymous):

65 mph is the average of 65 miles every hour

OpenStudy (anonymous):

ok so at 65mph, in 2 hours what is the mileage you were able to travel ??

OpenStudy (texaschic101):

Distance = Speed x Time (d = st) d = 395 s = 65 since we are looking for time, lets rearrange the equation... d = st d/s = t now lets sub d/s = t 395/65 = t 6.076 = t....you can probably round to 6.08 hours